OpenStudy (the_fizicx99):

-Must add up to 24 -Can only use set of given numbers \(\ once \) -Can only use \(\ basic \) operations, this includes multiplication, division, addition and subtract Given set: {1,3,4,6} In the booklet I was reading it said "a problem like this can be used to define how "hacking" is defined as." >.< How would you actually solve it, after trying 3 different methods, I re-read the instruction >.> I was off. My younger brother asked to use exponential equations. I didn't actually try, didn't seem to be much of use.

Parth (parthkohli):

Can we use brackets here? I came up with...\[6 \div (1 -(3\div 4))\]

Parth (parthkohli):

There's no multiplication in there though.

Parth (parthkohli):

Oh, no addition either.

hero (hero):

It's legal

OpenStudy (the_fizicx99):

It says you can use them, it's not necessary. And yes, that's one of the solutions.

OpenStudy (the_fizicx99):

There's one I'm really interested in solving, but it's midway through the book. They're scattered throughout the book. I read up to the 15th page just because I wanted to see if it matched my solution.

OpenStudy (the_fizicx99):

That would be 4 * 5 = 20 not 24

Parth (parthkohli):

Ooh, I found the one I wrote on Yahoo Answers too.

Parth (parthkohli):

Oh, dammit. That's actually 8.

OpenStudy (the_fizicx99):

How is that 8? o.o

Parth (parthkohli):

\[3+1 = 4\]\[6 -4 = 2\]

Parth (parthkohli):

Well, anyway, I hope the first answer really is legit.

OpenStudy (the_fizicx99):

Yes the first one is legit. There's a few more I believe. >.< I have to wait though .-.

terenzreignz (terenzreignz):

Forgive my naivete... but what's an exponential equation (in this context)?

Parth (parthkohli):

Probably exponents.

OpenStudy (the_fizicx99):

^ It involves logs @terenzreignz

Parth (parthkohli):

Ah lol, hacky way\[\log_3(1) + 4 \times 6\]

terenzreignz (terenzreignz):

I thought you meant actual exponents XD \[\Large 1^3(4\cdot6)\]

OpenStudy (the_fizicx99):

xD I thought about using them too. I wasn't allowed .-.

OpenStudy (the_fizicx99):

Well I have to get off for a while :/

Parth (parthkohli):

Whoa Terence, that was beautiful. ;_;

terenzreignz (terenzreignz):

You overreact parth XD Besides, exponents aren't allowed methinks?

OpenStudy (the_fizicx99):

Nope

Parth (parthkohli):

Yeah - but that solution was pretty cool.

OpenStudy (the_fizicx99):

Though exponents should be allowed >.> x^2 = x * x which is valid >.>

Parth (parthkohli):

But again, you can use something only once.

terenzreignz (terenzreignz):

It would involve using x twice, though, violating one of the rules ^_^

OpenStudy (the_fizicx99):

But that would be a repeating term, 1^3 = 1 * 1 * 1

OpenStudy (the_fizicx99):

Wow we all said that >.<

Parth (parthkohli):

lol

terenzreignz (terenzreignz):

Talk about 'repeating', eh :D

Parth (parthkohli):

ZOMG.

OpenStudy (anonymous):

..so who ever can solve this can say they are a hacker? ^_^

OpenStudy (anonymous):

\[\frac{ 6 }{ 1-(3/4) }\]

OpenStudy (the_fizicx99):

No, it said solving problems like this is what "defines" hacking. I'm still reading into it. The term hacking is tossed around allot.